OpenStudy (anonymous):

Multiply: 5 sqrt8 x4 sqrt7

OpenStudy (anonymous):

= 20sqrt56 =40sqrt14 you can use a calculator to work out what that is

OpenStudy (jdoe0001):

\(\bf 5\sqrt{8}\times 4\sqrt{7}\implies (5\times 4)\sqrt{8\times 7}\)

OpenStudy (anonymous):

im confused @jdoe0001

OpenStudy (jdoe0001):

just multiply coefficients and radical values

OpenStudy (anonymous):

still confused @jdoe0001 and @InspyR4D

OpenStudy (jdoe0001):

http://www.youtube.com/watch?v=AaDCAq1f-cg